OpenStudy (anonymous):

Which of the following is the sketch of the polynomial f(x) = 2(x - 3)(2x - 3)(2x + 1)(x + 2)?

OpenStudy (anonymous):

what are the choices?

OpenStudy (anonymous):

sketch will be shaped like a W, cross the x axis in four places: -2, -(1/2), 1.5 and 3

OpenStudy (anonymous):

the a and b is

OpenStudy (anonymous):

its A

OpenStudy (anonymous):

thanks and also i need help with this question? Which of the following is the sketch of the polynomial f(x) = x(x – 1)(x – 2)?

OpenStudy (anonymous):

the choices are

OpenStudy (anonymous):

if you could help sorry if i ask alot of questions. i took this last year so i dont remember.

OpenStudy (anonymous):

B

OpenStudy (anonymous):

and its cool

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

hey candy do you know how to Find the roots for the polynomial y = x(x + 1)(x – 3)(x – 2√2)(x + 2√2). i never heard of this when i took algebra.

OpenStudy (anonymous):

the choices are

OpenStudy (anonymous):

its D

OpenStudy (anonymous):

wait its not D

OpenStudy (anonymous):

Its B because when you solve for X on all of them you get X=0 x+1=0 -1 -1 x=-1 x-3=0 +3 +3 x=3

OpenStudy (anonymous):

and so on

OpenStudy (anonymous):

oh ok thanks i kinda understand now.

OpenStudy (anonymous):

np, ur welcome

OpenStudy (anonymous):

i meant to say she never taught use passed polynomials she never taught use parent graph.

OpenStudy (anonymous):

and i really need help with this question i went on youtube but the videos were no help i went on google but my computer didnt allow for me to view

OpenStudy (anonymous):

Which of the following is the sketch for g(x) = (x + 2)2 - 1 given the parent graph f(x) = (x)2?

OpenStudy (anonymous):